Find the inverse of the function. Determine whether the inverse is a function and state the domain and range.
f(x)=6-2x
1
Expert's answer
2011-10-14T12:41:07-0400
Find the inverse of the function. Determine whether the inverse is a function and state the domain and range. f(x)=6-2x

Let f(x)=y(x)

y(x)=6-2x => 2x=6-y(x) => x=3- y(x)/2=3-y/2=x(y)

Range of x(y)& (-inf; +inf)

Domain of y (-inf; +inf)